Q.4150 Two points on the demand curve for volleyballs are shown (P, Q) = (19, 55) and (P, Q) = (21, 45). Where P is price per volley ball in pounds and Q is Quantity demanded.

a) What is the elasticity of demand between these two points?

Answer _____________________________________________________________________


b) What is the point elasticity of demand at price 19 when price rises to 21?
1
Expert's answer
2020-09-01T11:09:59-0400

a) Answer: "\eta_{d} = -2"


Solution

The question requires arc elasticity of demand. Arc elasticity of demand calculates elasticity at the midpoint of the two given points; elasticity will not be affected by the direction of movement between the two points.


"\u2206P = |21 - 19|"

"= 2"


"\u2206Q = |45 - 55|"

"= 10"


Average price: "\hat P = \dfrac {19 + 21}{2}"


"= \dfrac {40}{2}"


"= 20"


Average quantity: "\hat Q = \dfrac {55 + 45}{2}"


"= \dfrac {100}{2}"


"= 50"


"\%∆P = \dfrac {∆P}{\hat P}×100\%"


"= \dfrac {2}{20}×100\%"


"= 10\%"



"\%∆Q = \dfrac {∆Q}{\hat Q}×100\%"


"= \dfrac {10}{50}×100\%"


"= 20\%"


Elasticity: "\eta_{d} = \dfrac {\%∆Q}{\%∆P}"


"= - \dfrac {20\%}{10\%}"


"= -2"



b) Answer: "\eta_{d} = -1.73"


Solution


"\u2206P = 21 - 19"

"= 2"



"\u2206Q = 45 - 55"

"= -10"



"\%∆P = \dfrac {∆P}{P_{0}}×100\%"


"= \dfrac {2}{19}×100\%"


"= 10.5263157\%"



"\%∆Q = \dfrac {∆Q}{Q_{0}}×100\%"


"= \dfrac {-10}{55}×100\%"


"= -18.1818181\%"



Elasticity: "\eta_{d} = \dfrac {\%∆Q}{\%∆P}"

"= \dfrac {-18.1818181\%}{+10.5263157\%}"


"= -1.72727273"


"= -1.73"